log(base x) 11 = 3
how do you solve this question

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

log(base x) 11 = 3
----
log(11)/log(x) = 3
----
log(x) = log(11)/3
---
log(x) = 3.1242
----
x = 10*3.1242
------
x = 1331.07
=================
Cheers,
Stan H.

Answer by Alan3354(69443)

log(x,11) = 3
x^3 = 11
x = cube root(11)
x =~ 2.22398